# Seeing Sarcasm Through Different Eyes: Analyzing Multimodal Sarcasm Perception in Large Vision-Language Models

### v2

@@ Abstract
With the advent of large vision-language models (LVLMs) demonstrating increasingly human-like abilities, a pivotal question emerges: do different LVLMs interpret multimodal sarcasm differently, and can a single model grasp sarcasm from multiple perspectives like humans? To explore this, we introduce an analytical framework using systematically designed prompts on existing multimodal sarcasm datasets. Evaluating 12 state-of-the-art LVLMs over 2,409 samples, we examine interpretive variations within and across models, focusing on confidence levels, alignment with dataset labels, and recognition of ambiguous “neutral” cases. Our findings reveal notable discrepancies—across LVLMs and within the same model under varied prompts. While classification-oriented prompts yield higher internal consistency, models diverge markedly when tasked with interpretive reasoning. These results challenge binary labeling paradigms by highlighting sarcasm’s subjectivity. We advocate moving beyond rigid annotation schemes toward multi-perspective, uncertainty-aware modeling, offering deeper insights into multimodal sarcasm comprehension. Our code and data are available at: https://github.com/CoderChen01/LVLMSarcasmAnalysis.
@@ I Introduction
Through comprehensive experiments on 12 cutting-edge LVLMs over 2,409 multimodal samples from the MMSD2.0 dataset [3], we observe notable variability among models as well as within individual models’ responses to different prompts. Notably, smaller models are more susceptible to prompt variations, whereas larger models display relatively stable performance yet still show biases toward sarcastic or literal interpretations. Moreover, explicit prompting for neutral judgments exposes considerable ambiguity beyond traditional binary labels. These findings highlight LVLMs’ potential for rich, human-like multimodal sarcasm comprehension while simultaneously revealing the constraints of current approaches, which rely heavily on datasets that reduce sarcasm to a rigid binary category. This calls for a paradigm shift toward uncertainty-aware, multi-perspective modeling methods that account for the subjective nuances of sarcasm.
@@ III-A Task Definition
[…]
Trinary Sarcasm Classification (TSC): To test whether LVLMs can directly identify neutral samples, we introduce the TSC task, defined as:
@@
[…]
@@ V-A1 Metric
As described in Section III-A, each task yields predicted classification labels, enabling a classification-based qualitative analysis. To quantify consistency, we use Krippendorff’s $\alpha$ [56] as the Classification Consistency Score, as it effectively handles partially missing or unavailable data—particularly relevant in cases where models fail to follow instructions, resulting in missing predictions for certain samples. However, a limitation of this metric is its potential instability in scenarios with extreme class distributions or small datasets. Krippendorff’s $\alpha$ has a range of $[-1,1]$, where a value closer to 1 indicates higher consistency, and a value closer to 0 indicates poorer consistency. Negative values generally indicate that the variance in the ratings between evaluators is greater than that expected by random consistency.
@@ V-A2 Result
Figure 4: Classification consistency heatmap. This figure illustrates the consistency scores of classification results from different prompt variants across various tasks for each model. Specifically, the gray blocks indicate that the Krippendorff’s $\alpha$ calculation for int2-8B in the BSC task is invalid. Further inspection reveals that int2-8B classifies all samples as sarcastic across all prompt variants, preventing the computation of the Krippendorff’s $\alpha$.
@@ VI-C Result
Aggregated results (Figure 7) show BSC aligns best with ground truth, while TSC shifts non-sarcastic samples toward neutrality. SCS and LCS results confirm that models are more confident in literal interpretations.
Overall, model agreement with ground truth varies. Sarcasm detection remains challenging, with a bias toward literal meaning. This highlights the limitations of binary-labeled datasets and the need for more nuanced evaluation frameworks. Detailed results are presented in Appendix A.
@@ VII-A Metric
For each task, we assess model confidence by aggregating the Negative Log-Likelihood (NLL) of predictions. A lower NLL (closer to 0) indicates higher confidence, whereas a higher NLL (closer to 1) signifies greater uncertainty.
@@ IX-B Inter-Model Subjectivity
Figure 12: This figure shows how different LVLMs give varying but reasonable interpretations when performing the BSC task on the same sample.
@@
X Conclusion
This study evaluated the ability of Large Vision-Language Models (LVLMs) to understand multimodal sarcasm using a novel framework comprising four tasks: Binary Sarcasm Classification (BSC), Ternary Sarcasm Classification (TSC), Sarcasm-Centric Scoring (SCS), and Literal-Centric Scoring (LCS). By analyzing 12 state-of-the-art LVLMs on 2,409 samples from MMSD2.0, we examined their consistency, alignment with dataset labels, and capacity to recognize neutral cases. Our findings show that LVLMs vary significantly in sarcasm interpretation, both across models and within the same model under different prompts. Larger models exhibit more stable performance but still lean toward either sarcastic or literal interpretations. The introduction of a neutral category in TSC highlights the limitations of binary-labeled datasets, as models frequently detect ambiguous cases. Additionally, LVLMs tend to favor literal interpretations unless explicitly prompted otherwise. These insights suggest the need to move beyond binary classification toward uncertainty-aware modeling. Future work should explore graded annotations and improved training strategies to enhance LVLMs’ ability to handle sarcasm in a more nuanced and flexible manner.

### v3

@@ Abstract
With the advent of large vision-language models (LVLMs) demonstrating increasingly human-like abilities, a pivotal question emerges: do different LVLMs interpret multimodal sarcasm differently, and can a single model grasp sarcasm from multiple perspectives like humans? To explore this, we introduce an analytical framework using systematically designed prompts on existing multimodal sarcasm datasets. Evaluating 12 state-of-the-art LVLMs over 2,409 samples, we examine interpretive variations within and across models, focusing on confidence levels, alignment with dataset labels, and recognition of ambiguous “neutral” cases. We further validate our findings on a diverse 100-sample mini-benchmark, incorporating multiple datasets, expanded prompt variants, and representative commercial LVLMs. Our findings reveal notable discrepancies—across LVLMs and within the same model under varied prompts. While classification-oriented prompts yield higher internal consistency, models diverge markedly when tasked with interpretive reasoning. These results challenge binary labeling paradigms by highlighting sarcasm’s subjectivity. We advocate moving beyond rigid annotation schemes toward multi-perspective, uncertainty-aware modeling, offering deeper insights into multimodal sarcasm comprehension. Our code and data are available at: https://github.com/CoderChen01/LVLMSarcasmAnalysis.
@@ I Introduction
Through comprehensive experiments on 12 cutting-edge LVLMs over 2,409 multimodal samples from the MMSD2.0 dataset [3], we observe notable variability among models as well as within individual models’ responses to different prompts. Notably, smaller models are more susceptible to prompt variations, whereas larger models display relatively stable performance yet still show biases toward sarcastic or literal interpretations. Moreover, explicit prompting for neutral judgments exposes considerable ambiguity beyond traditional binary labels. These findings highlight LVLMs’ potential for rich, human-like multimodal sarcasm comprehension while simultaneously revealing the constraints of current approaches, which rely heavily on datasets that reduce sarcasm to a rigid binary category. This calls for a paradigm shift toward uncertainty-aware, multi-perspective modeling methods that account for the subjective nuances of sarcasm. To test the robustness and generalizability of our conclusions, we further construct a diverse 100-sample mini-benchmark that integrates multiple datasets, expanded prompt variants, and representative commercial LVLMs. This supplementary analysis is complemented by human evaluation, which confirms that even state-of-the-art models produce plausible but interpretively variable outputs under different prompting conditions.
@@ III-A Task Definition
where $\mathcal{P}^{\textrm{BSC}}(\cdot)$ denotes the prompt template specifying the task, and $\hat{Y}^{\textrm{BSC}}_{i}\in\{\textrm{sarcastic},\textrm{non-sarcastic}\}$ represents the predicted label. Additionally, $R^{\textrm{BSC}}_{i}$ is the rationale for the prediction, and $S^{\textrm{BSC}}_{i}\in[0,1]$ is the confidence score. The inclusion of $R^{\textrm{BSC}}_{i}$ provides interpretability, allowing for a deeper analysis of the model’s reasoning process and potential biases in sarcasm detection.
Ternary Sarcasm Classification (TSC): To test whether LVLMs can directly identify neutral samples, we introduce the TSC task, defined as:
@@
where $\mathcal{P}^{\textrm{TSC}}(\cdot)$ is similar to $\mathcal{P}^{\textrm{BSC}}(\cdot)$, except that it enables the model to output ternary labels: $\hat{Y}^{\textrm{TSC}}_{i}\in\{\textrm{sarcastic},\textrm{non-sarcastic},\textrm{neutral}\}$. The inclusion of the neutral category enables a more nuanced assessment of sarcasm perception and ambiguity recognition in LVLMs.
@@ V-A1 Metric
As described in Section III-A, each task yields predicted classification labels, enabling a classification-based qualitative analysis. To quantify consistency, we use Krippendorff’s $\alpha$ [56] as the Classification Consistency Score, as it effectively handles partially missing or unavailable data—particularly relevant in cases where models fail to follow instructions, resulting in missing predictions for certain samples. However, a limitation of this metric is its potential instability in scenarios with extreme class distributions or small datasets. Krippendorff’s $\alpha$ has a range of [1,1][-1,1], where a value closer to 1 indicates higher consistency, and a value closer to 0 indicates poorer consistency. Negative values generally indicate that the variance in the ratings between evaluators is greater than that expected by random consistency.
@@ V-A2 Result
Figure 4: Classification consistency heatmap. This figure illustrates the consistency scores of classification results from different prompt variants across four tasks for each model. Specifically, the gray blocks indicate that the Krippendorff’s $\alpha$ calculation for int2-8B in the BSC task is invalid. Further inspection reveals that int2-8B classifies all samples as sarcastic across prompt variants, preventing the computation of Krippendorff’s $\alpha$.
@@ VI-C Result
As shown in Figure 7, aggregated results show BSC aligns best with ground truth, while TSC shifts non-sarcastic samples toward neutrality. SCS and LCS results confirm that models are more confident in literal interpretations.
Overall, model agreement with ground truth varies. Sarcasm detection remains challenging, with a bias toward literal meaning. This highlights the limitations of binary-labeled datasets and the need for more nuanced evaluation frameworks. Detailed results can be found in Appendix A.
@@ VII-A Metric
We assess model confidence using Negative Log-Likelihood (NLL), computed as the aggregated token-level log-probability of the model’s own output conditioned on the input (i.e., the image-text pair and prompt). Lower NLL values indicate higher model certainty, while higher values suggest greater uncertainty. NLL is a standard and theoretically grounded metric for autoregressive models, enabling consistent confidence estimation across both classification-oriented (BSC, TSC) and open-ended reasoning tasks (SCS, LCS). In the context of subjective phenomena like sarcasm, NLL serves as a proxy for the model’s interpretive commitment, indicating how decisively it favors one perspective under a given prompt. We emphasize that NLL is used as a within-model measure to analyze task- or prompt-level variations, rather than for direct comparison across architectures, which may differ in tokenization and decoding strategies. Overall, NLL offers an interpretable and model-native signal for assessing confidence in multimodal sarcasm interpretation.
@@ IX-B Inter-Model Subjectivity
Figure 12: An example illustrating inter-model subjectivity in sarcasm interpretation. Four LVLMs perform the BSC task on the same sarcastic text-image pair but yield divergent classifications with well-justified rationales. While lav1.5-7B and qw2-72B interpret the scene literally, citing intentional cosplay and absence of exaggerated tone, lav1.6-7B and int2-76B emphasize playful self-deprecation and cultural irony. This example demonstrates that different models prioritize different contextual cues (e.g., tone, attire, facial expression), supporting our claim that multimodal sarcasm is inherently subjective and that rigid binary labeling may fail to capture this diversity.
@@
X A Mini Mixed Dataset for Enhanced Evaluation
To complement our main evaluation and address key concerns regarding dataset diversity, prompt robustness, and model reliability, we conduct a focused small-scale experiment designed to test the generalizability and interpretive consistency of our framework under extended conditions. Specifically, we aim to investigate whether our core findings hold when applied to additional datasets with different stylistic and contextual features, a broadened set of prompt variants for each task, and representative commercial vision-language models not included in the main study.
Due to computational constraints and commercial API limitations, we restrict this supplementary analysis to a 100-sample mini-benchmark. Despite its limited scale, this targeted setup enables us to isolate and examine critical dimensions of model behavior, including prompt sensitivity, cross-dataset generalization, and the alignment of model outputs with human judgment, with improved granularity and interpretability.
X-A Dataset Construction
To assess the generalizability of our framework beyond a single benchmark, we construct a small-scale evaluation set by sampling from two distinct multimodal sarcasm datasets: MMSD2.0 [3] and DocMSU [39].
MMSD2.0 is a widely used benchmark for multimodal sarcasm detection, featuring carefully curated image-text pairs with binary sarcastic and non-sarcastic labels. Compared to its predecessor MMSD [2], MMSD2.0 mitigates spurious textual cues and label noise, making it a more reliable dataset for sarcasm analysis. DocMSU is a document-level benchmark designed to evaluate multimodal sarcasm understanding in longer and more context-rich samples. It includes sarcastic and non-sarcastic posts from Reddit  [^3] https://www.reddit.com, combining textual content with associated images or memes, and reflects more diverse linguistic and visual patterns. From each dataset, we randomly sample 25 sarcastic and 25 non-sarcastic instances, resulting in a balanced 100-sample mini-benchmark. This combined set captures stylistic variation across platforms, annotation strategies, and sarcasm manifestations, and provides a compact yet diverse basis for extended evaluation.
X-B Prompt Variant Scaling
While our main study uses three semantically equivalent prompt variants per task to evaluate inter-prompt consistency, some readers may reasonably question whether this number is sufficient to reveal the full extent of prompt sensitivity in LVLMs. After all, surface-level phrasing can introduce subtle but consequential variation in how models interpret the same task, particularly in subjective domains such as sarcasm [58, 47, 59]. To further explore this dimension, we expand each task to include ten prompt variants, all designed to be semantically equivalent but lexically diverse. This setup allows for a more rigorous examination of consistency under increased linguistic variation and provides a broader empirical basis for evaluating model robustness. We focus this experiment on two representative model families: Qwen2-VL [49] and InternVL2 [11]. These families were selected due to their strong performance in earlier experiments and their architectural contrast. Qwen2 models consistently demonstrated stable classification and rationale behavior, while InternVL2 models showed higher variability, making them ideal for studying prompt-induced instability.
As shown in Figure 13, we observe a general decline in classification consistency scores across tasks compared to the 3-prompt setting. Even in tasks like BSC and TSC, which are structurally more constrained, consistency drops for several models (e.g., int2-76B, qw2-7B). This suggests that prompt sensitivity is more substantial than previously estimated, highlighting that limited prompt variants may obscure important variation in model behavior. Similarly, rationale consistency (Figure 14) declines slightly across models and tasks, especially in SCS and LCS. Despite this, Qwen2 models retain a higher level of rationale stability than InternVL counterparts, echoing previous observations that model architecture plays a pivotal role in interpretive robustness. These findings confirm that prompt phrasing has non-trivial effects on sarcasm interpretation in LVLMs, and that increasing prompt diversity exposes more variation than smaller prompt sets reveal. We argue that future evaluations, especially for tasks involving subjectivity, should systematically account for prompt variation as a core evaluation axis rather than treating it as a minor implementation detail.
Figure 13: Classification consistency heatmap under expanded prompt settings.
Figure 14: Rationale consistency scores under extended prompt diversity.
X-C Evaluation on Commercial LVLMs
Figure 15: Quantitative analysis of commercial models.
To complement our analysis of open-source LVLMs and provide a more comprehensive picture of current multimodal sarcasm capabilities, we evaluate leading commercial models that represent the state-of-the-art in instruction-following and visual reasoning. While our primary experiments focus on publicly available models for reproducibility, commercial systems often demonstrate superior alignment and reasoning abilities, making them valuable baselines for this task. Specifically, we test six proprietary models: GPT-4o, GPT-4o-mini, GPT-4.1, GPT-4.1-mini, Gemini-2.5-Flash, and Gemini-2.5-Flash-lite. These models span different latency-accuracy trade-offs and are deployed via official APIs. All models are evaluated using the original three prompt variants per task on the 100-sample mini-benchmark introduced in Section X-A. This setup allows us to directly compare their behavior to open-source models under a consistent protocol.
We first analyze the inter-prompt consistency of commercial LVLMs across the four evaluation tasks. As shown in Figure 15(a) and (b), these models exhibit notably high agreement across semantically equivalent prompts. Krippendorff’s $\alpha$ scores remain positive and robust for all models, with near-perfect consistency observed in the BSC and TSC tasks for GPT-4o and GPT-4.1. This suggests that commercial systems possess strong surface-level prompt invariance, likely supported by more extensive instruction tuning and alignment strategies. However, despite this overall stability, a consistent drop is observed in the SCS and LCS tasks, where consistency is lower than in BSC and TSC across all models. This mirrors patterns seen in open-source models and points to a broader phenomenon: models, regardless of size or training, struggle more with interpretive tasks that require reasoning from a constrained perspective (e.g., sarcasm-only or literal-only), where subjective ambiguity cannot be resolved purely through prompt structure. One plausible explanation is that while commercial models are better aligned for general-purpose instruction following, they may default to a balanced interpretation strategy when faced with highly contextual or multi-intent stimuli, which undermines strict perspective adherence under prompt variation. Rationale consistency follows a similar trend: although overall coherence remains high, the gap between BSC/TSC and SCS/LCS again indicates that open-ended interpretive generation is inherently more sensitive to prompt phrasing, even for high-end models. These findings suggest that while alignment improves robustness to surface phrasing, interpretive rigidity under subjective constraints remains a fundamental challenge.
We then examine how commercial LVLMs compare with ground-truth labels, not as an absolute measure of performance, but to observe systematic patterns in model predictions. As shown in Figure 15(c) and (d), BSC task results are largely consistent with the annotated labels, with most errors arising from false positives, suggesting a slight tendency to over-assign sarcasm. In TSC, correctness decreases, especially for GPT models, mainly due to neutral classifications being excluded from evaluation, which highlights a mismatch between model uncertainty and rigid binary labels. SCS and LCS reveal distinct interpretive biases: sarcasm-centric prompting yields more false positives, while literal-centric prompting leads to sarcasm underprediction through false negatives. These trends persist in the COMP setting, where all models lean toward literal readings. These observations align with those found in open-source models, suggesting that even advanced commercial systems are shaped by the same label ambiguity and framing asymmetries inherent in sarcasm detection.
Finally, we analyze how commercial models identify neutral samples using the TSC and SCS-LCS strategies. As shown in Figure 15(e), commercial models detect relatively few neutral samples overall, with TSC consistently identifying fewer instances than SCS-LCS. This mirrors the behavior observed in open-source models and underscores the difficulty of eliciting ambiguity through direct classification. Among the models, GPT-4.1 and Gemini-2.5 Flash exhibit the highest consistency between the two methods, suggesting more stable neutral identification under different prompting conditions. However, for most models, the Jaccard overlap remains low, indicating that what is perceived as “neutral” still varies substantially across strategies even in aligned systems. Figure 15(f) further reveals that, similar to open-source models, neutral predictions in both methods are disproportionately drawn from non-sarcastic ground-truth samples. This bias suggests a general conservativeness in commercial models: when uncertain, they tend to interpret ambiguity as a weakened form of literalness rather than potential sarcasm. While this may reflect a rational default under annotation uncertainty, it also reinforces the limitations of using binary-labeled datasets for evaluating subjective phenomena like sarcasm, where ambiguity may legitimately arise on both sides of the label boundary.
In summary, commercial LVLMs show higher consistency and more stable behavior than open-source models, but still exhibit similar interpretive biases and difficulty with subjective prompts. We omit confidence-based analyses (e.g., NLL) in this section, as several commercial APIs do not provide log-probabilities, making direct comparison infeasible.
X-D Human Evaluation
Figure 16: UI interface of the human evaluation tool.
TABLE II: Distribution of Human Ratings Across Models and Tasks (%).
Rating BSC TSC SCS LCS
GPT-4o
Strong Disagr. (-3) 0.00 0.00 0.00 0.33
Mod. Disagr. (-2) 0.00 0.11 0.00 0.67
Disagreement (-1) 0.78 3.22 2.33 4.89
Uncertainty (0) 0.00 0.00 0.00 0.22
Agreement (+1) 27.67 35.44 32.22 35.00
Mod. Agreement (+2) 49.00 38.22 44.11 33.89
Strong Agreement (+3) 22.56 23.00 21.33 25.00
GPT-4.1
Strong Disagr. (-3) 0.00 0.11 0.11 0.11
Mod. Disagr. (-2) 0.11 0.11 0.33 0.78
Disagreement (-1) 1.00 1.44 4.11 5.22
Uncertainty (0) 0.00 0.00 0.00 0.11
Agreement (+1) 29.33 32.67 34.11 33.67
Mod. Agreement (+2) 47.11 39.33 37.44 34.89
Strong Agreement (+3) 22.44 26.33 23.89 25.22
Gemini-2.5-Flash
Strong Disagr. (-3) 0.00 0.00 0.22 0.00
Mod. Disagr. (-2) 0.11 0.00 0.56 0.44
Disagreement (-1) 0.89 0.78 7.33 3.33
Uncertainty (0) 0.11 0.11 0.00 0.11
Agreement (+1) 28.67 33.00 31.44 29.22
Mod. Agreement (+2) 46.22 41.78 36.78 37.78
Strong Agreement (+3) 24.00 24.33 23.67 29.11
TABLE III: Inter-Annotator Agreement (Krippendorff’s $\alpha$) across Tasks and Models.
Model BSC TSC SCS LCS
GPT-4o 0.14 0.07 0.07 0.14
GPT-4.1 0.19 -0.02 0.21 0.30
Gemini-2.5-Flash 0.29 -0.01 0.20 0.05
    Inter-annotator agreement (Krippendorff’s $\alpha$) is computed based on a three-level categorization of human ratings: agreement (scores +1, +2, +3), uncertainty (score 0), and disagreement (scores –1, –2, –3). This abstraction reduces Likert-scale granularity to focus on high-level consensus.
Given the inherently subjective nature of sarcasm, we complement our automated evaluation with a human assessment protocol to examine whether model outputs are subjectively plausible. Instead of asking annotators to perform the tasks themselves, we focus on validating the reasonableness of the model-generated predictions and rationales from a human perspective. To support this process, we design a custom human evaluation tool that allows annotators to view each multimodal input along with the model’s classification, score, and rationale for all four tasks. As shown in Figure 16, the interface presents the image, accompanying text, model metadata, and task-specific outputs, followed by a 7-point Likert-style agreement scale ranging from Strong Disagreement to Strong Agreement. Annotators evaluate each task independently, judging whether the given rationale sufficiently supports the predicted label. This setup enables us to assess interpretive plausibility without requiring full task replication by human raters.
To quantify human evaluation outcomes, we report two metrics for each model and task. First, we compute the distribution of Likert ratings to capture how plausible the model outputs appear to human judges. This provides insight into the subjective quality and perceived reliability of model reasoning under different prompt formulations. Second, we measure inter-annotator agreement using Krippendorff’s $\alpha$, which accommodates ordinal scales and handles missing or imbalanced data more robustly than alternative measures. This allows us to examine whether certain tasks or models produce more ambiguous or contentious predictions.
We apply this evaluation to the outputs of three representative commercial LVLMs: GPT-4o, GPT-4.1, and Gemini-2.5-Flash. For each model, we evaluate three prompt variants per task over the full 100-sample mini-benchmark introduced in Section X-A. Three annotators independently assess each output, with each annotator completing 3,600 judgments (3 models × 4 tasks × 3 prompt variants × 100 samples). This results in a total of 10,800 annotations across all annotators.
Our results reveal that the vast majority of model outputs were judged as subjectively plausible by human annotators. As shown in Table II, agreement ratings of 1 or higher (i.e., Agreement, Moderate Agreement, Strong Agreement) dominate across all models and tasks, particularly in the BSC setting. This indicates that, even without task-specific fine-tuning, commercial LVLMs can produce explanations that human raters find sensible and well-grounded.
From the perspective of inter-rater reliability, Table III shows that annotator agreement was consistently highest on the BSC task, with Krippendorff’s $\alpha$ scores of 0.14 (GPT-4o), 0.19 (GPT-4.1), and 0.29 (Gemini-2.5-Flash). This confirms that straightforward sarcasm classification is the most reliably judged by humans. In contrast, TSC yielded the lowest overall agreement, with values even falling into the negative range for GPT-4.1 and Gemini. This reflects the challenge of interpreting neutrality consistently. SCS and LCS also show distinctive trends across models. In SCS, GPT-4.1 achieved relatively high inter-annotator agreement ($\alpha=0.21$), while GPT-4o’s $\alpha=0.07$ was the lowest, despite both models receiving largely positive human ratings. This suggests that although GPT-4o’s rationales were generally perceived as agreeable (Table II), their framing may have been more variable or ambiguous, leading to less consistent interpretation across annotators. For LCS, Gemini-2.5-Flash showed weaker agreement ($\alpha=0.05$), despite strong Likert scores. This implies that annotators may disagree not about the quality of the rationale itself, but about whether it truly embodies a literal perspective.
Together, these results suggest that commercial LVLMs generate outputs broadly acceptable to humans, especially for direct classification tasks. However, constrained interpretive tasks like SCS and LCS introduce greater subjectivity and variability, even when output plausibility remains high. This underscores the importance of multi-perspective evaluation for understanding model behavior beyond label alignment.
XI Discussion
Our study demonstrates that large vision-language models (LVLMs), when evaluated through a carefully designed multi-perspective framework, reveal substantial interpretive variability in multimodal sarcasm understanding. These variations occur not only across model architectures, but also within the same model under different prompt phrasings and task formulations. This reinforces our central argument: sarcasm is inherently subjective and context-sensitive, and current binary-labeled datasets do not fully capture this complexity.
Through extensive experiments across 12 open-source models and a supplementary benchmark involving commercial systems and human evaluators, we identify two consistent findings: (1) LVLMs display stronger confidence and agreement on literal interpretations than sarcastic ones, and (2) interpretive prompts (e.g., SCS and LCS) introduce greater response variability than classification prompts (e.g., BSC and TSC), even in high-performing models. These findings point to a need for both uncertainty-aware modeling and more flexible annotation schemes that acknowledge interpretive ambiguity.
At the same time, we recognize that our current framework adopts a strictly zero-shot prompting strategy to ensure uniform evaluation across models. While this allows us to isolate intrinsic model behavior, it may also limit the expressive reasoning capabilities of LVLMs, particularly in tasks requiring constrained interpretive perspectives. Prior work such as MemeGuard [60] and VIGNETTE [61] has shown that prompting strategies like few-shot demonstrations, question decomposition, or chain-of-thought reasoning can improve model reliability in subjective or socially grounded tasks. Incorporating such techniques into our framework could mitigate prompt sensitivity, enhance rationale consistency, and better reflect human-like reasoning. We view these methods as promising next steps. For example, few-shot prompting may help models internalize contextual contrast more effectively, while chain-of-thought prompts may encourage models to explicitly differentiate between literal surface meaning and ironic intent. Additionally, socially informed prompting strategies, as proposed in VIGNETTE, could offer templates for disambiguating sarcasm from humor or criticism through guided sub-questioning. Due to resource constraints and our focus on establishing a zero-shot benchmark, we leave the empirical validation of these methods to future work, but emphasize their conceptual alignment with the goals of our framework.
Overall, our findings provide a strong baseline for evaluating subjective multimodal phenomena with LVLMs. They also lay the groundwork for richer prompting strategies and more cognitively aligned modeling approaches that go beyond label agreement toward interpretive plausibility and human-consistent understanding.
XII Conclusion
This study introduced a four-task evaluation framework to assess how Large Vision-Language Models (LVLMs) interpret multimodal sarcasm from multiple perspectives. Evaluating 12 open-source LVLMs on MMSD2.0, we found that models vary widely in their sarcasm judgments, with larger models showing more stability but still exhibiting interpretive bias. The TSC task revealed that many samples fall between sarcastic and non-sarcastic, highlighting the limitations of binary labels. Models also tended to favor literal interpretations unless explicitly guided otherwise. To test generalizability, we conducted a supplementary experiment on a 100-sample mini-benchmark combining MMSD2.0 and DocMSU, using more diverse prompts and including commercial LVLMs like GPT-4o. The results confirmed our main findings, and a human evaluation further showed that most model outputs were perceived as plausible, though variability increased under interpretive constraints. Overall, our findings call for sarcasm evaluation frameworks that move beyond binary classification toward uncertainty-aware, multi-perspective modeling that better reflects human interpretive diversity.